When U 238 nucleus originally at rest, decays by emitting an alpha particle having a speed u, the recoil speed of the residual nucleus is ?

Grade:12

1 Answers

Anmol gupta
5 years ago
By conservation of linear momentum. Mass of left uranium atom =238-4=234 as it has decayed an alpha particle whose mass is 4amu. 4×U=234×v where v is speed of nucleus     So v=4U÷234 is required answer